OpenStudy (anonymous):

please solve A=a+b+c/3, for c

OpenStudy (anonymous):

3A-a-b do all you people go to the same school or something? I saw this same question asked once before today

OpenStudy (anonymous):

heyy

OpenStudy (anonymous):

thats...actually incorrect. A-a-b=c/3 3(A-a-b)=c 3A-3a-3b=c

OpenStudy (anonymous):

i think they meant A=(a+b+c)/3